# Wandering around: A bioinspired approach to visual attention through object motion sensitivity

### v1

@@ Abstract
The system, characterised using ideal gratings and benchmarked against the Event Camera Motion Segmentation Dataset (EVIMO), reaches a mean IoU of 82.2% and a mean SSIM of 96% in multi-object motion segmentation. Additionally, the detection of salient objects reaches an accuracy of 88.8% in office scenarios and 89.8% in challenging indoor and outdoor low-light conditions, as evaluated on the Event-Assisted Low-Light Video Object Segmentation Dataset (LLE-VOS). A real-time demonstrator showcases the system’s capabilities of detecting the salient object through object motion sensitivity in $0.124$ s in dynamic scenes. Its learning-free design ensures robustness across diverse perceptual scenes, making it a reliable foundation for real-time robotic applications and serving as a basis for more complex architectures.
@@ 1 Introduction
Active vision entails dynamically adjusting camera viewpoints to enhance environmental perception and information gathering. The concept gained prominence in the late 1980s through the work of Aloimonos et al., who demonstrated that an active observer can solve basic vision problems more effectively than a passive one [aloimonos1987active]. In contrast, passive perception has often been described as challenging because it fails to resolve ill-posed problems [hadamard1902problemes] (existence, uniqueness and stability), lacking a unique solution and constraints [poggio1984analog, aloimonos1987active]. An active approach addresses the limitations of fixed-camera systems, such as occlusions and restricted fields of view. By adapting dynamically to changing conditions, active vision enables a more comprehensive and unobstructed visual experience, something vital for navigating the physical world. A key challenge in real-time robotics applications is balancing power consumption, latency, and data storage. Although traditional computer vision methods have achieved remarkable performance, these approaches struggle with generalisation, particularly in handling variations in scale, pose, and lighting conditions, or when identifying objects in complex or obstructed environments [zhu2020review]. Increasing data usage to overcome these challenges often exacerbates issues related to data processing, storage and transfer [feng2019computer].
@@
Given the agent’s finite computational abilities, it is essential to allocate limited perceptual and cognitive resources. Attention mechanisms are frequently employed in robotics [azizevaluation, pasquale2016enabling, aarthi2020study] to selectively process Regions of Interest (ROIs), effectively reducing the volume of data to be processed and thereby facilitating real-time responsiveness. Visual attention in robotics offers numerous benefits, such as enhancing interaction efficiency by enabling robots to selectively focus on relevant stimuli consequently improving adaptability and responsiveness in dynamic environments [hanifi2024pipeline, potapova2017survey, pasquale2016enabling].
@@
The process of filtering relevant information from the continuous bombardment of complex sensory data and shifting the center of the visual field (fovea) toward the current ROI, mimics the biological mechanism of overt selective attention [rizzolatti1983mechanisms]. This process occurs not only in animals, where selecting the most significant stimuli (such as the presence of a predator) is crucial but also in complex machinery with a diverse array of sensors, such as robots, where similar attentional mechanisms can be leveraged to focus the processing on the most salient part of the visual field. Models of bottom-up visual attention [itti1998model, itti2001computational, ude2005distributed] have been proposed to generate feature-based saliency maps for detecting salient regions or proto-objects [russell2014model].
@@
To detect objects, the human visual perception of a scene strongly depends on perceptual organisation rules, known as the ’Gestalt laws’ [kohler1967gestalt], such as continuity, proximity, figure-ground organisation, and closure. These principles give perceptual saliency to regions of the visual field that can be perceived as “proto-objects”, referring to areas of the scene that can potentially contain an object. To detect the presence of a proto-object, perceptual grouping mechanisms of the visual input refer to neuronal populations of the secondary visual cortex (V2) [zhou2000coding] encoding the geometric properties of a proto-object through their firing patterns. The sensitivity of these cells extends beyond edge orientation to encompass the concept of “border ownership”, defining the ownership of the edges to foreground or background finally detecting proto-objects.
@@
Several frame-based proto-object approaches have utilised a bioinspired architecture resembling border ownership cells [russell2014model] by applying the von Mises kernel with varying orientations and scales [todorovic2008gestalt] to identify areas of the visual field that potentially contain objects. These implementations feed different cues of information into the proto-object model, generating saliency maps dependent on the provided input, such as motion [molin2013proto], depth [hu2016proto] and texture [uejima2020proto, uejima2018proto]. These models have significantly influenced our understanding of vision and have inspired robotic implementations [orabona2005object].
@@
However, these models predominantly rely on traditional frame-based vision sensors, which are limited by restricted dynamic range, redundancy, motion blur, latency and high power consumption. Event-based cameras [Lichtsteiner2008], which mimic the initial layers of the mammalian retina and react to pixel-level illumination changes, offer a solution to these limitations. Unlike frame-based cameras, event-based sensors improve dynamic range, reduce latency, and generate an asynchronous event stream providing information of spatial coordinates, polarity, and timestamps. This results in a significant reduction in data processing, making event-based cameras highly relevant for robotic applications [monforte2020exploiting, mueggler2018continuous, Iacono2018, glover2017robust]. Their inherent real-time response to luminance changes provides an ideal sensory input for guiding subsequent visual attention actions. Previous implementations of event-driven proto-object models for the humanoid robot iCub [bartolozzi2011embedded], leveraging cues of intensity [iacono2019proto] and depth [ghosh2022event], have paved the way for fully bioinspired proto-object models in robotics, effectively bridging the gap between bioinspired algorithms and hardware.
@@
While these models utilise the sparse, event-driven nature of visual sensors, their full potential is unrealised without execution on a neuromorphic platform. Most computer vision processing occurs on CPUs, GPUs, or specialized devices like Tensor Processing Units (TPUs) and Neural Processing Units (NPUs), which are optimized for dense data but require constant data transfer between the sensor, memory, and processor [MutluGhoseGomez-LunaEtAl_2019_ProcessingDataWhere], consuming vast amounts of energy. Frame-based cameras also consume significant energy by capturing data continuously. Neuromorphic computing [douglas1995neuromorphic] addresses these challenges by leveraging event-driven sensing and Spiking Neural Networks (SNNs) to process streams of events in parallel, inspired by the brain’s efficiency [DiehlS2015SpikingDeepNetworks, 7838165, gehrig2020eventbasedangularvelocityregression, SpikeMS] demonstrating to be effective also for robotics real-time applications [fatahi2024event, Dangelo_etal22, d2020event]. Unlike von Neumann architectures, neuromorphic systems co-locate computation and memory, reducing data transfer and achieving greater energy efficiency. By mimicking biological neurons, neuromorphic circuits, whether analog [pehle2022brainscales, moradi2017scalable, qiao2015reconfigurable] or digital [Furber2014, davies2018loihi, sga2020neuromorphic], overcome traditional processing limitations through SNNs and integrated memory and computation units. A further advancement was the implementation of the proto-object model proposed by Iacono et al. [iacono2019proto] on the SpiNNaker neuromorphic platform [Dangelo_etal22]. This effort [Dangelo_etal22] led to the development of the first end-to-end, saliency-based proto-object Spiking Neural Network (SNN), significantly reducing processing latency from $\sim$100 ms to approximately $\sim$16 ms.
@@
All of the mentioned event-based and neuromorphic proto-object models implemented on the humanoid robot iCub [iacono2019proto, ghosh2022event, Dangelo_etal22] rely on stereotypical circular eye movements that mimic human fixational eye movements to perceive static scenes. The sparse event-based output from the sensors generates events not only in response to changes in the visual scene but also due to egomotion-induced motion of the camera. These egomotion-induced events overlay the relative motion of objects, requiring a mechanism to enhance the detection of objects’ motion. Object Motion Sensitivity (OMS) is critical to distinguishing between self-motion and objects’ motion in the visual field, which, in turn, is essential for shifting attention towards moving items (such as prays and predators), and for effective navigation and interaction with the environment [schwartz2021object].
@@
Recent advancements in object motion sensitivity models have provided valuable insight into how biological visual systems enhance object motion during eye movements. Ölveczky et al. [olveczky2003segregation] and later Baccus et al. [baccus2008retinal] investigated the neural circuitry underlying OMS in retinal ganglion cells, revealing their ability to selectively respond to relative motion between the center and the surrounding regions of the receptive fields. These cells suppress responses to global image motion, such as that generated by an observer’s movements, enabling the identification of moving objects against dynamic backgrounds. These studies further pinpointed the involvement of specific retinal interneurons, including fast, transient bipolar cells and polyaxonal amacrine cells, which mediate the OMS response through carefully balanced excitatory and inhibitory pathways.
@@
These cells have inspired recent advancements in modelling object motion sensitivity systems, deriving biologically motivated circuits [schwartz2021object]. The hardware implementation, Integrated Retinal Functionality in Image Sensors (IRIS) [yin2023iris, sinaga2024hardwarealgorithmreengineeringretinalcircuit] proposed a CMOS-compatible design of retinal-like computations, such as OMS, to enable low-power real-time in-sensor processing with minimal computational overhead. Further work by Clerico et al. [clerico2024retina] proposed an algorithmic implementation of OMS showing state-of-the-art performance [parameshwara2021spikems, parameshwara20210, zhang2023multi] while utilising fewer parameters, making it particularly suitable for resource-constrained systems. By processing Dynamic Vision Sensor (DVS) data, the algorithm achieves a significant reduction in computational complexity by a factor of 1000 compared to traditional methods [parameshwara2021spikems, parameshwara20210, zhang2023multi], making it suitable for low-latency, low-bandwidth applications in dynamic environments. The process employed by these algorithms to enhance relative motion does not ensure the identification of whether the motion results from a coherently moving object in the visual field or from a disjointed set of features that do not belong to a single object (i.e., clutter), thereby necessitating a mechanism for perceptual grouping such as the saliency-based proto-object model [Dangelo_etal22].
@@
The proposed work builds upon the visual attention work by D’Angelo et al.[Dangelo_etal22, ghosh2022event, iacono2019proto] and the object motion sensitivity model by Clerico et al.[clerico2024retina], combining them to reduce the amount of information processed by focusing solely on the salient proto-object of the scene while enhancing object motion during camera motion. In an effort to develop an architecture that significantly reduces power consumption and latency, this work bridges the gap between bio-inspired software and hardware by proposing a fully spiking Convolutional Neural Network (sCNN) architecture that leverages event-based sensors and neuromorphic algorithms. Furthermore, the learning-free design of both algorithms ensures their applicability across diverse scenarios, demonstrating the system’s capability to operate in new environments without requiring training datasets.
@@
The online architecture, which enables the system to wander around, enhance object motion detection through fixational eye movements, and perform saccades toward salient objects, is implemented using a Speck device mounted on a Pan-Tilt Unit (PTU). This setup explores the environment by processing events generated by the integrated DVS sensor on the Speck. The Speck [^2] https://www.synsense.ai/products/speck-2/ platform is the world’s first fully event-driven neuromorphic vision System-on-Chip (SoC) integrating a state-of-the-art dynamic vision sensor (DVS) with a fully asynchronous chip architecture. In the pursuit of fast online robotic applications, opting for the Speck neuromorphic hardware from the array of available hardware solutions emerges as the optimal choice due to its seamless integration with a DVS demonstrating low energy consumption [sorbaro2020optimizing] and suitability for common robotics online tasks [liu2019live]. The events from the DVS sensor are processed through the Spiking Object Motion Sensitivity (sOMS) model to enhance motion segmentation and through the SNN Proto-object model to focus visual attention on a single maximum salient proto-object. The identified maximum salient point is subsequently fed into the Spiking Attention Control (sAC) model to move the PTU and position the maximum salient point at the center of the visual field. The system then executes small stabilisation movements, similar to fixational eye movements in the mammalian visual system [ko2010microsaccades], to identify the next salient point and close the loop. Such small and frequent movements are essential to prevent the rapid fading of static elements in the environment, which are crucial for building a saliency map for the next saccade. An overview of the architecture is shown in Figure 1.
@@
  • Characterisation of the Spiking Object Motion Sensitivity CNN model (sOMS) through qualitative and quantitative assessments, inspired by the cell recordings conducted by Ölveczky et al. [olveczky2003segregation], and an investigation of the parameters in the model proposed by Clerico et al. [clerico2024retina].
  • Characterisation of the SNN Proto-Object model investigating the model’s parameters, as demonstrated in [iacono2019proto, ghosh2022event, Dangelo_etal22], using the SalMapIROS dataset from [iacono2019proto].
  • Qualitative and Quantitative assessment of the sOMS model demonstrating the suppression of meaningless events and benchmarking the model against the EVIMO dataset [mitrokhin2019ev] computing Intersection over Union (IoU) and Structural Similarity Index (SSIM) scores.
  • Qualitative and quantitative benchmark to highlight the importance of incorporating the SNN Proto-object model for multi-object detection on the EVIMO dataset [mitrokhin2019ev].
  • Qualitative and quantitative benchmark in challenging low-light indoor and outdoor scenarios on the LLE-VOS dataset [mitrokhin2019ev] to demonstrate the system robustness.
@@ 2 Related work
Active vision enables intelligent systems to interact dynamically with their environment through mechanisms like selective visual attention, focusing on relevant details to achieve efficient and adaptive perception [rizzolatti1983mechanisms]. The human visual system is specifically structured to manage the continuous influx of stimuli from the environment. In the retina, receptive fields decrease non-linearly in size from the peripheral retina toward the fovea [harvey2011relationship]. Consequently, in many mammalian visual systems, the fovea captures a small, highly detailed portion of the scene, while the peripheral retina gathers the remainder at a lower resolution [Biologyfoveation]. Peripheral vision identifies points of interest and directs the fovea to focus on them sequentially; a process known as foveation. This mechanism balances wide-area scanning with a detailed inspection, enhancing efficiency and enabling the recognition of items within the scene [DigitalFoveation].
@@
Foveation is based on the principle that different regions of the visual field are processed at varying resolutions. Numerous systems have been proposed to emulate the log-polar mapping and spatially variant sampling distribution characteristic of the visual field [chessa2016space, GrimesDickeyPishko_2006_EncyclopediaSensors]. This subsampling of visual input offers several advantages, including data reduction [d2020event], motion direction detection [d2020event], enhanced attention mechanisms, improved object tracking and segmentation, and egomotion segmentation [BolducLevine_1998_ReviewBiologicallyMotivated]. The properties of foveated vision have been studied through both software-based simulation and hardware implementations. For instance, as first demonstrated by Geisler et al. [geisler1998real] in 1998 and later confirmed by De Figueiredo et al. [de2023overview], foveal vision systems are beneficial for reducing computational demand while maintaining effective performance in complex tasks, thereby providing foundational insight into space-variant, non-uniform vision mechanisms. In this regard, Dauce et al. [Dauce_2018_ActiveFoveaBasedVision] have demonstrated that highly accurate active inference can be achieved by applying saccadic movements using only a fraction of the raw visual input by applying a deep neural network as a predictive system. Deep learning model can also achieve both high throughput and high accuracy for saccadic attention with a log-polar foveated sampling of the input [LukanovKonigPipa_2021_BiologicallyInspiredDeep]. An active system with a foveated sensor with log-polar pixel mapping for tracking objects in the visual field [YamamotoYeshurunLevine_1996_ActiveFoveatedVision] was found to closely replicate the fixation behaviour of humans scanning objects in the visual field. In [LamTsangMengEtAl_2006_NeuromorphicTranslationalEgomotion], an FPGA-driven hardware platform for egomotion estimation is implemented on a mobile robotic system, demonstrating that log-polar sampling of the visual input can be harnessed to extract the optical flow and the heading direction of the robot. Comprehensive reviews of foveated systems for robotics can be found in [BolducLevine_1998_ReviewBiologicallyMotivated, JavierTraverBernardino_2010_ReviewLogpolarImaging].
@@
Selective attention in computer vision can also be approximated by representing the majority of a scene at a low resolution while maintaining high detail in specific key regions, ROIs. This approach selectively discards less relevant information, preserving essential details to enhance task-specific performance without implementing any log-polar mapping. To that end, a common approach involves using pairs of sensors to emulate the effects of convergence and disparity estimation [ChessaMurgiaNardelliEtAl_2014_BioinspiredActiveVision]. For instance, a foveated stereo-matching algorithm [xu2016evaluation] can significantly improve the efficiency of robotic cloth manipulation tasks, such as grasping and flattening, by achieving comparable accuracy at two to three times the speed of traditional methods. In [Scassellati_1999_BinocularFoveatedActive], a binocular foveated system is implemented on a robotic head that can simulate saccadic motion and smooth pursuit, albeit with a significant cumulative pixel error. Additionally, Medeiros et al. [medeiros2020dynamic] introduced a dynamic multifoveated structure that allows robots to track and focus attention on multiple objects simultaneously without redundancy in processing, thereby optimizing real-time image analysis. This technique enhances visual attention capabilities in robotic systems by enabling rapid focus shifts across different regions of interest. Nevertheless, despite the advantages of visual processing with coupled sensors, such as superior object segmentation, considerable effort has been invested in achieving similar results with a single sensor due to the additional complexity necessary for synchronisation and parallel processing when using multiple sensors.
@@
Recent advancements in selective attention systems for neuromorphic vision have highlighted their potential to enhance visual processing efficiency in robotics due to their low power consumption and low latency. A pivotal work [gruel2023stakes] introduces neuromorphic foveation, using event cameras to selectively process visual information, reducing data load while maintaining high accuracy in tasks like semantic segmentation and classification. This approach achieves a superior balance between data quantity and quality compared to traditional high or low-resolution event data. While foveated vision enhances certain aspects of perception, it introduces an inherent challenge in event processing: distinguishing between independent object motion (i.e., object motion) and camera-induced motion (i.e., egomotion) while executing fixational eye movements and identifying the next salient point for foveation. Neuromorphic cameras detect intensity changes, making them sensitive to both objects and egomotion-induced events, which can create ambiguity. Disambiguating visual stimuli and perceptually grouping the information to focus the attention requires mechanisms to enhance object motion cues akin to biological vision systems, to ensure reliable object detection and tracking.
@@
Risi et al.[9948687] introduced an event-based FPGA implementation of a saliency-driven selective attention model that efficiently processes visual information from event cameras. This implementation outperforms classic bottom-up models, such as those by Itti et al.[itti1998model, itti2001computational], in predicting eye fixations, achieving high performance with minimal computational resources. It captures significant aspects of visual perception by focusing on localised temporal changes and provides microsecond-precise saliency updates without relying on traditional image frames. Event-based visual attention for robotic applications has already been proposed [rea2013event], harnessing its responsiveness and robustness under varying lighting conditions while mimicking human-like attentional processes.

### v2

@@ Abstract
The system, characterised using ideal gratings and benchmarked against the Event Camera Motion Segmentation Dataset (EVIMO), reaches a mean IoU of 82.2% and a mean SSIM of 96% in multi-object motion segmentation. Additionally, the detection of salient objects reaches an accuracy of 88.8% in office scenarios and 89.8% in challenging indoor and outdoor low-light conditions, as evaluated on the Event-Assisted Low-Light Video Object Segmentation Dataset (LLE-VOS). A real-time demonstrator showcases the system’s capabilities of detecting the salient object through object motion sensitivity in $0.124$ seconds in dynamic scenes. Its learning-free design ensures robustness across diverse perceptual scenes, making it a reliable foundation for real-time robotic applications and serving as a basis for more complex architectures.
@@ 1 Introduction
Active vision entails dynamically adjusting camera viewpoints to enhance environmental perception and information gathering. The concept gained prominence in the late 1980s through the work of Aloimonos et al. [1], who demonstrated that an active observer can solve basic vision problems more effectively than a passive one. In contrast, passive perception has often been described as challenging because it fails to resolve ill-posed problems [2] (such as existence, uniqueness, and stability), lacking both a unique solution and constraints [3, 1]. This limitation results in ambiguities, making it difficult to determine a single, definitive interpretation of the visual data. An active approach addresses the limitations of fixed-camera systems, such as occlusions and restricted fields of view. By adapting dynamically to changing conditions, active vision enables a more comprehensive and unobstructed visual experience, something vital for navigating the physical world. A key challenge in real-time robotics applications is balancing power consumption, latency, and data storage. Although traditional computer vision methods have achieved remarkable performance, these approaches struggle with generalisation, particularly in handling variations in scale, pose, and lighting conditions, or when identifying objects in complex or obstructed environments [4]. Increasing data usage to overcome these challenges often exacerbates issues related to data processing, storage and transfer [5].
@@
Given the agent’s finite computational abilities, it is essential to allocate limited perceptual and cognitive resources. Attention mechanisms are frequently employed in robotics [6, 7, 8] to selectively process Regions of Interest (ROIs), effectively reducing the volume of data to be processed and thereby facilitating real-time responsiveness. Visual attention in robotics offers numerous benefits, such as enhancing interaction efficiency by enabling robots to selectively focus on relevant stimuli consequently improving adaptability and responsiveness in dynamic environments [9, 10, 7].
@@
The process of filtering relevant information from the continuous bombardment of complex sensory data and shifting the centre of the visual field (fovea) toward the current ROI, mimics the biological mechanism of overt selective attention [11]. This process occurs not only in animals, where selecting the most significant stimuli (such as the presence of a predator) is crucial but also in complex machinery with a diverse array of sensors, such as robots, where similar attentional mechanisms can be leveraged to focus the processing on the most salient part of the visual field. Models of bottom-up visual attention [12, 13, 14] have been proposed to generate feature-based saliency maps for detecting salient regions or proto-objects [15].
@@
To detect objects, the human visual perception of a scene strongly depends on perceptual organisation rules, known as the ’Gestalt laws’ [16], such as continuity, proximity, figure-ground organisation, and closure. These principles give perceptual saliency to regions of the visual field that can be perceived as “proto-objects”, referring to areas of the scene that can potentially contain an object. To detect the presence of a proto-object, perceptual grouping mechanisms of the visual input refer to neuronal populations of the secondary visual cortex (V2) [17] encoding the geometric properties of a proto-object through their firing patterns. The sensitivity of these cells extends beyond edge orientation to encompass the concept of “border ownership”, which defines the assignment of edges to the foreground or background, ultimately enabling the detection of proto-objects.
@@
Several frame-based proto-object approaches have utilised a bioinspired architecture resembling border ownership cells [15] by applying the von Mises kernel with varying orientations and scales [18] to identify areas of the visual field that potentially contain objects. These implementations feed different cues of information into the proto-object model, generating saliency maps dependent on the provided input, such as motion [19], depth [20] and texture [21, 22]. These models have significantly influenced our understanding of vision and have inspired robotic implementations [23].
@@
However, these models predominantly rely on traditional frame-based vision sensors, which are limited by restricted dynamic range, redundancy, motion blur, latency and high power consumption. Event-based cameras [24], which mimic the initial layers of the mammalian retina and react to pixel-level illumination changes, offer a solution to these limitations. Unlike frame-based cameras, event-based sensors improve dynamic range, reduce latency, and generate an asynchronous event stream providing information of spatial coordinates, polarity, and timestamps. This results in a significant reduction in data processing, making event-based cameras highly relevant for robotic applications [25, 26, 27, 28]. Their inherent real-time response to luminance changes provides an ideal sensory input for guiding subsequent visual attention actions. Previous implementations of event-driven proto-object models for the humanoid robot iCub [29], leveraging cues of intensity [30] and depth [31], have paved the way for fully bioinspired proto-object models in robotics, effectively bridging the gap between bioinspired algorithms and hardware.
@@
While these models utilise the sparse, event-driven nature of visual sensors, their full potential is unrealised without execution on a neuromorphic platform. Most computer vision processing occurs on CPUs, GPUs, or specialized devices like Tensor Processing Units (TPUs) and Neural Processing Units (NPUs), which are optimized for dense data but require constant data transfer between the sensor, memory, and processor [32], consuming vast amounts of energy. Frame-based cameras also consume significant energy by capturing data continuously. Neuromorphic computing [33] addresses these challenges by leveraging event-driven sensing and Spiking Neural Networks (SNNs) to process streams of events in parallel, inspired by the brain’s efficiency [34, 35, 36, 37] demonstrating to be effective also for robotics real-time applications [38, 39, 40]. Unlike von Neumann architectures, neuromorphic systems co-locate computation and memory, reducing data transfer and achieving greater energy efficiency. By mimicking biological neurons, neuromorphic circuits, whether analog [41, 42, 43] or digital [44, 45, 46], overcome traditional processing limitations through SNNs and integrated memory and computation units. A further advancement was the implementation of the proto-object model proposed by Iacono et al. [30] on the SpiNNaker neuromorphic platform [39]. This effort [39] led to the development of the first end-to-end, saliency-based proto-object Spiking Neural Network (SNN), significantly reducing processing latency from $\sim$100 ms to approximately $\sim$16 ms.
@@
All of the mentioned event-based and neuromorphic proto-object models implemented on the humanoid robot iCub [30, 31, 39] rely on stereotypical circular eye movements that mimic human fixational eye movements to perceive static scenes. The sparse event-based output from the sensors generates events not only in response to changes in the visual scene but also due to egomotion-induced motion of the camera. These egomotion-induced events overlay the relative motion of objects, requiring a mechanism to enhance the detection of objects’ motion. Object Motion Sensitivity (OMS) is critical to distinguishing between self-motion and objects’ motion in the visual field, which, in turn, is essential for shifting attention towards moving items (such as prey and predators), and for effective navigation and interaction with the environment [47].
@@
Recent advancements in object motion sensitivity models have provided valuable insight into how biological visual systems enhance object motion during eye movements. Ölveczky et al. [48] and later Baccus et al. [49] investigated the neural circuitry underlying OMS in retinal ganglion cells, revealing their ability to selectively respond to relative motion between the center and the surrounding regions of the receptive fields. These cells suppress responses to global image motion, such as that generated by an observer’s movements, enabling the identification of moving objects against dynamic backgrounds. These studies further pinpointed the involvement of specific retinal interneurons, including fast, transient bipolar cells and polyaxonal amacrine cells, which mediate the OMS response through carefully balanced excitatory and inhibitory pathways.
@@
These cells have inspired recent advancements in modelling object motion sensitivity systems, deriving biologically motivated circuits [47]. The hardware implementation, Integrated Retinal Functionality in Image Sensors (IRIS) [50, 51] proposed a CMOS-compatible design of retinal-like computations, such as OMS, to enable low-power real-time in-sensor processing with minimal computational overhead. Further work by Clerico et al. [52] proposed an algorithmic implementation of OMS showing state-of-the-art performance [53, 54, 55] while utilising fewer parameters, making it particularly suitable for resource-constrained systems. By processing DVS data, the algorithm achieves a significant reduction in computational complexity by a factor of 1000 compared to traditional methods [53, 54, 55], making it suitable for low-latency, low-bandwidth applications in dynamic environments. The process employed by these algorithms to enhance relative motion does not ensure the identification of whether the motion results from a coherently moving object in the visual field or from a disjointed set of features that do not belong to a single object (i.e., clutter), thereby necessitating a mechanism for perceptual grouping such as the saliency-based proto-object model [39].
@@
The proposed work builds upon the visual attention work by D’Angelo et al.[39, 31, 30] and the object motion sensitivity model by Clerico et al.[52], combining them to reduce the amount of information processed by focusing solely on the salient proto-object of the scene while enhancing object motion during camera motion. In an effort to develop an architecture that significantly reduces power consumption and latency, this work bridges the gap between bio-inspired software and hardware by proposing a fully spiking Convolutional Neural Network (sCNN) architecture that leverages event-based sensors and neuromorphic algorithms. Furthermore, the learning-free design of both algorithms ensures their applicability across diverse scenarios, demonstrating the system’s capability to operate in new environments without requiring training datasets.
@@
The online architecture, which enables the system to wander around, enhance object motion detection through fixational eye movements, and perform saccades toward salient objects, is implemented using a Speck device mounted on a Pan-Tilt Unit (PTU). This setup explores the environment by processing events generated by the integrated DVS sensor on the Speck. The Speck [^2] https://www.synsense.ai/products/speck-2/ platform is the world’s first fully event-driven neuromorphic vision System-on-Chip (SoC) integrating a state-of-the-art DVS with a fully asynchronous chip architecture. In the pursuit of fast online robotic applications, opting for the Speck neuromorphic hardware from the array of available hardware solutions emerges as the optimal choice due to its seamless integration with a DVS demonstrating low energy consumption [56] and suitability for common robotics online tasks [57]. The events from the DVS sensor are processed through the Spiking Object Motion Sensitivity (sOMS) model to enhance motion segmentation and through the SNN Proto-object model to focus visual attention on a single maximum salient proto-object. The identified maximum salient point is subsequently fed into the Spiking Attention Control (sAC) model to move the PTU and position the maximum salient point at the center of the visual field. The system then executes small stabilisation movements, similar to fixational eye movements in the mammalian visual system [58], to identify the next salient point and close the loop. Such small and frequent movements are essential to prevent the rapid fading of static elements in the environment, which are crucial for building a saliency map for the next saccade. An overview of the architecture is shown in Figure 1.
@@
  • Characterisation of the Spiking Object Motion Sensitivity CNN model (sOMS) through qualitative and quantitative assessments, inspired by the cell recordings conducted by Ölveczky et al. [48], and an investigation of the parameters in the model proposed by Clerico et al. [52].
  • Characterisation of the SNN Proto-Object model investigating the model’s parameters, as demonstrated in [30, 31, 39], using the SalMapIROS dataset from [30].
  • Qualitative and Quantitative assessment of the sOMS model demonstrating the suppression of meaningless events and benchmarking the model against the EVIMO dataset [59] computing Intersection over Union (IoU) and Structural Similarity Index (SSIM) scores.
  • Qualitative and quantitative benchmark to highlight the importance of incorporating the SNN Proto-object model for multi-object detection on the EVIMO dataset [59].
  • Qualitative and quantitative benchmark in challenging low-light indoor and outdoor scenarios on the LLE-VOS dataset [60] to demonstrate the system robustness.
@@ 2 Related work
Active vision enables intelligent systems to interact dynamically with their environment through mechanisms like selective visual attention, focusing on relevant details to achieve efficient and adaptive perception [11]. The human visual system is specifically structured to manage the continuous influx of stimuli from the environment. In the retina, receptive fields decrease non-linearly in size from the peripheral retina toward the fovea [61]. Consequently, in many mammalian visual systems, the fovea captures a small, highly detailed portion of the scene, while the peripheral retina gathers the remainder at a lower resolution [62]. Peripheral vision identifies points of interest and directs the fovea to focus on them sequentially; a process known as foveation. This mechanism balances wide-area scanning with a detailed inspection, enhancing efficiency and enabling the recognition of items within the scene [63].
@@
Numerous systems have been proposed to emulate the log-polar mapping and spatially variant sampling distribution characteristic of the visual field [64, 65]. This subsampling of visual input offers several advantages, including data reduction [40], motion direction detection [40], enhanced attention mechanisms, improved object tracking and segmentation, and egomotion segmentation [66]. The properties of foveated vision have been studied through both software-based simulation and hardware implementations. For instance, as first demonstrated by Geisler et al. [67] in 1998 and later confirmed by De Figueiredo et al. [68], foveal vision systems are beneficial for reducing computational demand while maintaining effective performance in complex tasks, thereby providing foundational insight into space-variant, non-uniform vision mechanisms. In this regard, Dauce et al. [69] have demonstrated that highly accurate active inference can be achieved by applying saccadic movements using only a fraction of the raw visual input by applying a deep neural network as a predictive system. Deep learning models can also achieve both high throughput and high accuracy for saccadic attention through log-polar foveated sampling of the input [70]. An active system with a foveated sensor with log-polar pixel mapping for tracking objects in the visual field [71] was found to closely replicate the fixation behaviour of humans scanning objects in the visual field. In [72], an FPGA-driven hardware platform for egomotion estimation is implemented on a mobile robotic system, demonstrating that log-polar sampling of the visual input can be harnessed to extract the optical flow and the heading direction of the robot. Comprehensive reviews of foveated systems for robotics can be found in [66, 73].
@@
Selective attention in computer vision can be approximated by representing the majority of a scene at low resolution while preserving high detail in key regions of interest (ROIs), without implementing any log-polar mapping. To that end, a common approach involves using pairs of sensors to emulate the effects of convergence and disparity estimation [74]. For instance, a foveated stereo-matching algorithm [75] can significantly improve the efficiency of robotic cloth manipulation tasks, such as grasping and flattening, by achieving comparable accuracy at two to three times the speed of traditional methods. In [76], a binocular foveated system is implemented on a robotic head that can simulate saccadic motion and smooth pursuit, albeit with a significant cumulative pixel error. Additionally, Medeiros et al. [77] introduced a dynamic multifoveated structure that allows robots to track and focus attention on multiple objects simultaneously without redundancy in processing, thereby optimizing real-time image analysis. This technique enhances visual attention capabilities in robotic systems by enabling rapid focus shifts across different regions of interest. Nevertheless, despite the advantages of visual processing with coupled sensors, such as superior object segmentation, considerable effort has been invested in achieving similar results with a single sensor due to the additional complexity necessary for synchronisation and parallel processing when using multiple sensors.
@@
Recent advancements in selective attention systems for neuromorphic vision have highlighted their potential to enhance visual processing efficiency in robotics due to their low power consumption and low latency. A pivotal work [78] introduces neuromorphic foveation, using event cameras to selectively process visual information, reducing data load while maintaining high accuracy in tasks like semantic segmentation and classification. This approach achieves a superior balance between data quantity and quality compared to traditional high or low-resolution event data. While foveated vision enhances certain aspects of perception, it introduces an inherent challenge in event processing: distinguishing between independent object motion (i.e., object motion) and camera-induced motion (i.e., egomotion) while executing fixational eye movements and identifying the next salient point for foveation. Disambiguating visual stimuli and perceptually grouping the information to focus the attention requires mechanisms to enhance object motion cues akin to biological vision systems, to ensure reliable object detection and tracking.
@@
Risi et al.[79] introduced an event-based FPGA implementation of a saliency-driven selective attention model that efficiently processes visual information from event cameras. This implementation outperforms classic bottom-up models, such as those by Itti et al.[12, 13], in predicting eye fixations, achieving high performance with minimal computational resources. It captures significant aspects of visual perception by focusing on localised temporal changes and provides microsecond-precise saliency updates without relying on traditional image frames. Event-based visual attention for robotic applications has already been proposed [80], harnessing its responsiveness and robustness under varying lighting conditions while mimicking human-like attentional processes.